OpenStudy (anonymous):

solve the system x+3y=1 -2x+y=-23 PLEASE HELP

OpenStudy (abb0t):

Do you know substitution method? oR elimination?

OpenStudy (anonymous):

NO

OpenStudy (abb0t):

Well, first, you want to rearrange one of the equations above into the form x=f(y) or y=f(x)

OpenStudy (abb0t):

That means rearrange the first one x = 1-3y then plug that new "x" into the second equation, combine like terms, and solve for \(y\)